MONT 107Q – Thinking About Mathematics Second Writing Assignment March 8, 2017

For the second of the larger paper assignments this semester, you have the choice either

• To look in detail at one of the stories in Ovid, and then trace its “afterlife” in popular culture. This last part will require some research and so you will need to follow usual procedures for a research paper, document your sources carefully, etc. or

• To write a metamorphosis of your own in prose or poetry along the lines of one of Ovid’s stories, together with an explanation for what you were trying to do.

The expectations and requirements for each of these options are slightly different, so please read the following carefully and make note of what you are asked to do before starting to work(!) Please submit via Google Docs or as a MS Word .doc or .docx file to the usual address: jlittle@holycross.edu by 5:00pm on Monday, March 20.

Topic 1 – Pygmalion

One of the most famous of the stories in the Metamorphoses is told in a relatively brief section of Book X (pages 335 - 338 in the Mandelbaum translation). The tale of the sculptor Pygmalion and his creation of “a female figure more exquisite than a woman who was born could ever match” has many twists and turns. By implication at least, it also has a happy ending (despite several actively “creepy” aspects of what Pygmalion does and despite the fact that we might expect a disaster on the grounds that he has usurped the prerogatives of the gods by making such a life-like statue to start with!) For this paper topic, you should start out by giving your analysis of Ovid’s version of the story. How and why are the metamorphosis in this story different from those in others we read? In particular, you should consider the implications of the fact that the goddess Venus apparently looks with favor on the marriage of Pygmalion and his creation and grants them a normal married life together, children, etc. What does this story say about the roles of art and love in human life? Is the Pygmalion story about love? about art? about something else?

If you do some digging, you will see that this story from Ovid (and other forms of the Greek myth that Ovid drew on) has inspired an extraordinary number of other works of art (paintings, sculptures, plays, etc.) from the 1600’s on. Two in particular are the play Pygmalion by George Bernard Shaw, and the Broadway musical and film My Fair Lady based on the Shaw play. Pick one of these and explain: How does it use aspects of the Ovid story? How does it introduce new elements? What does your example say about the influence of Ovid on popular culture? Expectations and Requirements: For this topic, your paper should be roughly 5 pages (double-spaced, one side)–you may find that you have more to say than that; “go for it” if you do(!) As on the other writing assignments, pay careful attention to your opening paragraph, make your transitions clear and easy to follow for the reader, and aim for no technical errors. Include a bibliography of all the sources you consulted and use foot- or end-notes for any direct quotations you use or statements of fact that are not legitimate “general knowledge.” Online sources are OK, but please place the full URL and the date you accessed the page for each online source in your bibliography.

Topic 2 – A Metamorphosis of One’s Own(!)

Part a. Imagine that you are Ovid transported to 2017 and something has survived the transition: your desire to sing of metamorphoses(!) For this paper topic, you will write your own metamorphosis story. Don’t feel constrained by the actual language or poetic forms that you (Ovid) used to use, though – speak in your own voice as it is now. Begin by choosing a figure (modern, mythological, literary, or historical) and think about how and why it would be transformed into someone or something else. Following what you have seen of Ovid’s style in the Metamorphoses, write a 300-400 word Metamorphosis story. Feel free to get creative and have fun concocting a tale of transformation. You could metamorphose Hilary Clinton into a panda, a panda into the actor Jack Black, or Jack Black into an intelligent computer operating system like the one played by Scarlett Johansson in the 2013 film Her. (Don’t use any of these ideas – come up with one of your own!!) But try to stay as close as possible to Ovid’s (i.e. your!) methods.

Part b. Discuss your story. Include a 100-200 word commentary examining your newly written Metamorphosis. In your discussion, answer some or all of the questions below. If any of the original Metamorphoses inspired you, identify which these were.

1. Exactly where does the metamorphosis occur within your story? Does it begin the story or serve as its climax? What effects does this metamorphosis have on the story and the character?

2. What form does the metamorphosis take?

3. Why does this metamorphosis occur?

4. Which of the following best describes your attitude toward the metamorphosis? Is it tragic, grotesque, comic, etc.? How does the way you told the story reflect that attitude?

5. Post-metamorphosis, is any part of the character left unchanged? How or why does that part not change? Is the unchanged part left that way to torment another character? Is it a comforting reminder of the past for the character who was changed? Note: This commentary is required in addition to the 300-400 word story.

Expectations and Requirements: No sources other than the Mandelbaum translation of Ovid, please!